# A Prospect-Theoretic Policy Gradient Algorithm for Behavioral Alignment in Reinforcement Learning

### v2

@@ G.2 Proof of Proposition 2
G.3 Proof of Proposition 3
We build a classical MDP which is equivalent to our CPT-MDP following the procedure described in more details in appendix G.1. In summary, we expand every state $s$ in our CPT MDP into several states encoding the state, partial sum of rewards and timestep all at once, i.e. $\tilde{s}=(s,t,\sum_{k=0}^{t-1}r_{k})$. We encode all the dynamics of our CPT MDP into our new conventional MDP, and the reward in our new MDP when going from a state $\tilde{s}_{a}=(s_{a},\sigma_{a},t_{a})$ to another state $\tilde{s}_{b}=(s_{b},\sigma_{b},t_{a}+1)$ is simply defined as $\mathcal{U}(\sigma_{b})-\mathcal{U}(\sigma_{a})$. We see that any non-Markovian policy in the CPT MDP can be rewritten as a non-Markovian policy in the new classical MDP and reciprocally, any non-Markovian policy in the new, classical MDP can be rewritten as a non-Markovian policy in the original CPT MDP. Moreover, choosing $1$ as a discounting factor in the new MDP, we note that by telescoping sum, the total reward in the new MDP corresponds exactly to the total utility of the sum of rewards in the CPT MDP. Since it is a classical MDP, our new MDP admits a Markovian optimal policy [Puterman, 2014]. This optimal policy is at least as good as any other non-Markovian policy in the new MDP with regards to the expected total reward and therefore at least as good as any other non-Markovian policy in the old CPT MDP with regards to expected utility of the total reward. When translating this optimal policy to a policy in the CPT MDP, we notice it only depends on $(s,\sum_{k=0}^{t-1}r_{k},t)$, meaning it is indeed an element of $\Pi_{\Sigma,NS}$. This concludes the proof.
G.4 Proof of Theorem 4
We will prove the following extended version of Theorem 4.
Theorem 12.
Let $\mathcal{U}$ be continuous and (strictly) increasing. The following statements are equivalent: 1. For any MDP, there exists an optimal policy for (EUT-PO) in $\Pi_{M,NS}$ 2. There exists a function $\varphi:\mathbb{R}^{2}\to\mathbb{R}$ such that: $\forall x,a,b\in\mathbb{R},b\neq 0,\mathcal{U}(x+a)-\mathcal{U}(x)=\varphi(a,b)(\mathcal{U}(x+b)-\mathcal{U}(x))\,.$ 3. There exists $\alpha\in\mathbb{R}$ s.t. $\mathcal{U}^{\prime\prime}(x)=\alpha\,\mathcal{U}^{\prime}(x)$ for all $x\in\mathbb{R}\,.$ 4. There exist $A,B,C\in\mathbb{R}$ s.t. $\mathcal{U}(x)=Ax+B$ or $\mathcal{U}(x)=A+B\exp(Cx)$ for all $x\in\mathbb{R}\,.$ 5. There exists a function $\mu:\mathbb{R}^{2}\to\mathbb{R}$ such that: $\forall y,c,d\in\mathbb{R},\mathcal{U}(y+c)-\mathcal{U}(c)=\mu(c,d)(\mathcal{U}(y+d)-\mathcal{U}(d))\,.$
We prove a series of implications and equivalences. It can be easily verified from combining all these results that $1\implies 2\implies 3\implies 4\implies 5\implies 1\,,$ which proves all the equivalences of the theorem. We proceed to prove each one of our implications in the rest of this section.
Proof of $3\Leftrightarrow 4$, $4\Rightarrow 2$, and $4\Rightarrow 5$. The equivalence $3\Leftrightarrow 4$ is obtained simply by solving the differential equation for one implication and a simple calculation for the other implication. The implications $4\Rightarrow 2$ and $4\Rightarrow 5$ follow from simple algebraic verification.
Proof of $5\Rightarrow 2$.
We suppose 5 holds. For any given $a,b\in\mathbb{R}$ such that $b\neq 0$, we define $\varphi(a,b):=\frac{\mathcal{U}(1+a)-\mathcal{U}(1)}{\mathcal{U}(1+b)-\mathcal{U}(1)}\,.$ Notice that this quantity is well defined since $\mathcal{U}$ being (strictly) increasing (and $b\neq 0$) implies that $\mathcal{U}(1+b)-\mathcal{U}(1)\neq 0\,.$ Then, we use 5 to obtain that for every $x\in\mathbb{R},$
$\mathcal{U}(x+b)-\mathcal{U}(x)=\mu(x,1)(\mathcal{U}(1+b)-\mathcal{U}(1))\,.$ (11)
We conclude the proof of the implication by writing:
$\mathcal{U}(x+a)-\mathcal{U}(x)$ $=\mu(x,1)(\mathcal{U}(1+a)-\mathcal{U}(1))$ (again by 5)
$=\varphi(a,b)\mu(x,1)(\mathcal{U}(1+b)-\mathcal{U}(1))$ (using the above definition of $\varphi(a,b)$)
$=\varphi(a,b)(\mathcal{U}(x+b)-\mathcal{U}(x))\,.$ (using Eq. (11))
This shows that 2 holds and concludes the proof.
Proof of $2\Rightarrow 4$. Consider a fixed integer $k$. Let $C_{k}:=\varphi(2\cdot 2^{-k},2^{-k})$ and $u_{n}:=\mathcal{U}(n2^{-k})$ for every $n\in\mathbb{N}$. We have the following recurrence relation for all $n\in\mathbb{N}$:
$u_{n+2}-u_{n}=C_{k}(u_{n+1}-u_{n})\,.$
That is to say for every $n\in\mathbb{N}$,
$u_{n+2}-C_{k}u_{n+1}+(C_{k}-1)u_{n}=0\,.$
This recurrence relation can be solved by examining the characteristic polynomial:
$x^{2}-C_{k}x+(C_{k}-1)=0\,.$
The roots are obtained using the quadratic formula: $r_{\pm}=\frac{C_{k}\pm\sqrt{C_{k}^{2}-4(C_{k}-1)}}{2}=\frac{C_{k}\pm(C_{k}-2)}{2}\,.$
  • If $C_{k}=2$, there is only one root, $1$, and $\exists\,D_{k},E_{k}\in\mathbb{R},\forall n,u_{n}=(D_{k}n+E_{k})1^{n}=D_{k}n+E_{k}$
  • Otherwise, there are two real roots, $(C_{k}-1)$ and $1$, and $u_{n}$ is of the form, $u_{n}=D_{k}(C_{k}-1)^{n}+E_{k}\cdot 1^{n}=D_{k}C_{k}^{n}+E_{k}\,.$
This proves that for all $k$, there exists a function in $\{x\mapsto Ax+b,(A,B)\in\mathbb{R}^{2}\}\bigcup\{x\mapsto A+B\exp(Cx),(A,B)\in\mathbb{R}^{2}\}$ that coincides with $\mathcal{U}$ on the set $\{\frac{x}{2^{k}},x\in\mathbb{N}\}$.
Importantly, all these functions have to be the same (across different values of $k$, i.e. all $D_{k}$ constants are the same and all $E_{k}$ constants do also coincide), due to the structure of $\{x\mapsto Ax+b,(A,B)\in\mathbb{R}^{2}\}\bigcup\{x\mapsto A+B\exp(Cx),(A,B)\in\mathbb{R}^{2}\}$ and because they all coincide on all the integers with the corresponding value of the same (fixed) utility function $\mathcal{U}$ at the relevant integer. This means that there exists a single function $f$ in $\{x\mapsto Ax+b,(A,B)\in\mathbb{R}^{2}\}\bigcup\{x\mapsto A+B\exp(Cx),(A,B)\in\mathbb{R}^{2}\}$ which coincides with $\mathcal{U}$ on all of $\{\frac{x}{2^{y}},x\in\mathbb{N},y\in\mathbb{N}^{+}\}$. By continuity of $\mathcal{U}$, we obtain that 4 holds.
Proof of $4\Rightarrow 1$.
If $\mathcal{U}$ is an affine function $x\mapsto Ax+B$ (for some $A,B\in\mathbb{R}$), then solving the (EUT-PO) problem boils down to solving a traditional MDP in which an optimal Markovian policy always exists [Puterman, 2014].
Let us now assume that the utility function is of the form $x\mapsto A+B\exp(Cx)$ for some $A,B,C\in\mathbb{R}$. Without loss of generality, we can simply ignore the constant $A$ in the optimization problem (EUT-PO) and just assume we are maximizing $\mathcal{U}(x)=B\exp(Cx)$. Recall that we are considering a finite-horizon setting with horizon length $H$. For any $0\leq T\leq H$, we say that a policy $\pi\in\Pi_{\Sigma,NS}$ is Markovian in the last $T$ steps if there exists a function $f$ defined from $\mathcal{S}\times\mathbb{N}$ (into the set of policies) such that:
$\forall\sigma\in\mathbb{R},\forall t\geq H-T,\forall s\in\mathcal{S},\pi(s,\sigma,t)=f(s,t)\,.$
Using again the unwinded MDP construction like in the proof of Proposition 3, we can find a policy $\pi^{\star}\in\Pi_{\Sigma,NS}^{D}$ which is ”totally” optimal: that is to say, starting from any $(s,\sigma,t)$, $\mathbb{E}\left[\mathcal{U}(\sigma+\sum_{k=t}^{H-1}r_{k})\right]$ is maximal when following policy $\pi^{\star}$. We proceed by induction to prove the assertion $\mathcal{P}_{T}$: ‘There exists a deterministic totally optimal policy $\pi_{T}$ which is Markovian in the last $T$ steps’ for any $T\leq H$, especially for $T=H$ which is the desired result.
Initialization: $\mathcal{P}_{0}$ is true with $\pi_{0}=\pi^{\star}$.
Induction: Let us suppose $\mathcal{P}_{T}$ is true for some $T<H$. We define $\pi_{T+1}$ by:
$\pi_{T+1}(s,\sigma,t):=\begin{cases}\pi_{T}(s,0,t)&\text{if }t=H-T-1\\ \pi_{T}(s,\sigma,t)&\text{otherwise}\,.\end{cases}$
We see that $\pi_{T+1}$ is a deterministic policy that is Markovian in the last $T+1$ steps. We also see that for any $t\geq H-T$ and any $\sigma\in\mathbb{R},s\in\mathcal{S}$, starting from $(s,\sigma,t)$, $\mathbb{E}\left[\mathcal{U}(\sigma+\sum_{k=t}^{H-1}r_{k})\right]$ is maximal when following policy $\pi_{T+1}$. We need to prove it for others values of $t$. i.e. $t\leq H-T-1$.
Because $\pi_{T+1}$ is Markovian in the last $T+1$ steps, the probability distribution on future states, actions and rewards starting from $(s,\sigma,H-T-1)$ does not depend on $\sigma$.
We know that it optimizes $\mathbb{E}\left[\mathcal{U}(0+\sum_{k=t}^{H-1}r_{k})\right]$, and we want to show that it optimizes $\mathbb{E}\left[\mathcal{U}(\sigma+\sum_{k=H-T-1}^{H-1}r_{k})\right]$ for all $\sigma\in\mathbb{R}$. This is where we use the form of the utility function to remark that
$\mathbb{E}\left[\mathcal{U}\left(\sigma+\sum_{k=H-T-1}^{H-1}r_{k}\right)\right]=\mathbb{E}\left[\exp(C\sigma)\mathcal{U}\left(\sum_{k=H-T-1}^{H-1}r_{k}\right)\right]=\exp(C\sigma)\,\mathbb{E}\left[\mathcal{U}\left(\sum_{k=H-T-1}^{H-1}r_{k}\right)\right]$
and a maximizer for $\sigma=0$ is therefore a maximizer for all $\sigma$.
We know now that $\pi_{T+1}$ is optimal starting from any $(s,\sigma,t)$ if $t\geq H-T-1$. Note that now, step $H-T-1$ is included.
Starting from $(s,\sigma,t)$ with $t<H-T-1$, we know that $\pi_{T}$ maximizes $\mathbb{E}\left[\mathcal{U}(\sigma+\sum_{k=t}^{H-1}r_{k})\right]$. We notice, using the tower rule:
$\mathbb{E}\left[\mathcal{U}\left(\sigma+\sum_{k=t}^{H-1}r_{k}\right)\right]$ $=\mathbb{E}\left[\mathcal{U}\left(\sigma+\sum_{k=t}^{H-T-2}r_{k}+\sum_{k=H-T-1}^{H-1}r_{k}\right)\right]$
$=\mathbb{E}\left[\mathbb{E}\left[\mathcal{U}\left(\underbrace{\sigma+\sum_{k=t}^{H-T-2}r_{k}}_{\sigma^{\prime}}+\sum_{k=H-T-1}^{H-1}r_{k}\right)|\underbrace{\sigma+\sum_{k=t}^{H-T-2}r_{k}}_{\sigma^{\prime}},s_{H-T-1}\right)\right]\,.$
Because $\pi_{T+1}$ is as good as $\pi_{T}$ at maximizing $\mathbb{E}\left[\mathcal{U}\left(\sigma^{\prime}+\sum_{k=H-T-1}^{H-1}\gamma^{k}r_{k}\right)\right]$ starting from $(\sigma^{\prime},S_{H-T-1},T-H-1)$, we conclude that $\pi_{T+1}$ performs as well as $\pi_{T}$, because the inner conditional expectation is the same and the first steps are the same.
Therefore, $\mathcal{P}_{T+1}$ is true.
Conclusion: $\mathcal{P}_{H}$ is true, which is our desired result.
Proof of $1\Rightarrow 2$.
Let us show $\neg 2\Rightarrow\neg 1$. $\neg 2$ means that for any function $\varphi:\mathbb{R}^{2}\mapsto\mathbb{R}$, there exists $x,a,b\in\mathbb{R}$ such that $b\neq 0$ and $\mathcal{U}(x+a)-\mathcal{U}(x)\neq\varphi(a,b)(\mathcal{U}(x+b)-\mathcal{U}(x))\,.$
Define now $\varphi:\mathbb{R}^{2}\mapsto\mathbb{R}$ by $\varphi(\alpha,\beta)=\frac{\mathcal{U}(\alpha)-\mathcal{U}(0)}{\mathcal{U}(\beta)-\mathcal{U}(0)}$ for all $\alpha\in\mathbb{R},\beta\neq 0$ and $\varphi(\alpha,\beta)=1$ for $\beta=0$. It follows that there exist $x,a\in\mathbb{R},b\neq 0$ (given by $\neg 2$ above) such that $\mathcal{U}(x+a)-\mathcal{U}(x)\neq\varphi(a,b)(\mathcal{U}(x+b)-\mathcal{U}(x))\,.$
As a consequence, we obtain $\frac{\mathcal{U}(x+a)-\mathcal{U}(x)}{\mathcal{U}(x+b)-\mathcal{U}(x)}\neq\frac{\mathcal{U}(a)-\mathcal{U}(0)}{\mathcal{U}(b)-\mathcal{U}(0)}\,.$ Our idea now is to exploit this difference in utility to build a situation in which a non-Markovian strategy is clearly more profitable in view of our (EUT-PO) policy optimization problem.
Suppose without loss of generality that $b>a>0$ and $\frac{\mathcal{U}(x+a)-\mathcal{U}(x)}{\mathcal{U}(x+b)-\mathcal{U}(x)}>\frac{\mathcal{U}(a)-\mathcal{U}(0)}{\mathcal{U}(b)-\mathcal{U}(0)}$ without loss of generality. In the other cases, the inequalities might get reversed but the gist of the proof stays the same. We define $p$ as the halfpoint
$p:=\frac{1}{2}\left(\frac{\mathcal{U}(x+a)-\mathcal{U}(x)}{\mathcal{U}(x+b)-\mathcal{U}(x)}+\frac{\mathcal{U}(a)-\mathcal{U}(0)}{\mathcal{U}(b)-\mathcal{U}(0)}\right)\,.$ (12)
Since $\mathcal{U}$ is strictly increasing, we have that $p\in(0,1)\,.$
We now consider an MDP with three states $s_{0},s_{1},s_{2}$ where $s_{2}$ is a terminal state that leads to nowhere and $s_{0}$ is the starting state. Whatever action is taken in $s_{0}$, we transition to $s_{1}$, with reward $x$ with probability $\frac{1}{2}$ and reward $y$ with probability $\frac{1}{2}$. Once in state $s_{1}$, we can take action A, which yields reward $a$ with probability $1$, or take action $B$, which yields reward $b$ with probaility $p$ and 00 otherwise. Both lead to $s_{2}$ and the end of the episode with certainty. Here, to maximize the (EUT-PO) objective, one has to adopt a non-Markovian strategy in $s_{1}$, hence disproving assertion $1$. Indeed, knowing the reward achieved in the past step (between states $s_{0}$ and $s_{1}$) allows to decide whether to take more risks or not to achieve a higher EUT return.
We elaborate on this claim in what follows. Observe first that
$\frac{\mathcal{U}(x+a)-\mathcal{U}(x)}{\mathcal{U}(x+b)-\mathcal{U}(x)}>p>\frac{\mathcal{U}(a)-\mathcal{U}(0)}{\mathcal{U}(b)-\mathcal{U}(0)}\,.$ (13)
What is the best action to choose if we are in state $s_{1}$ and have had return 00 so far? We know there is a deterministic best action to take. Action $1$ yields total reward $\mathcal{U}(a)$. Choosing action $2$ yields:
$\mathbb{E}(\mathcal{U}(r_{0}+r_{1})|r_{0}=0)=\mathbb{E}(\mathcal{U}(r_{1}))=p\,\mathcal{U}(b)+(1-p)\,\mathcal{U}(0)=p(\mathcal{U}(b)-\mathcal{U}(0))+\mathcal{U}(0)>\mathcal{U}(a)\,,$ (14)
where the strict inequality follows from using (13). So it is strictly better to choose action $2$ over action $1$ if we are in $s_{1}$ and have had a return 00 so far.
What is the best action to choose if we are in state $s_{1}$ and have had return $x$ so far? We know again that there is a deterministic best action to take. Action $1$ yields total reward $\mathcal{U}(x+a)$. Choosing action $2$ yields:
$\mathbb{E}(\mathcal{U}(r_{0}+r_{1})|r_{0}=x)=\mathbb{E}(\mathcal{U}(r_{1}+x))=p\,\mathcal{U}(b+x)+(1-p)\,\mathcal{U}(x)\\ =p(\mathcal{U}(b+x)-\mathcal{U}(x))+\mathcal{U}(x)<\mathcal{U}(a+x)\,,$ (15)
where the strict inequality follows from using again (13). So it is strictly better to choose action $1$ over action $2$ if we are in $s_{1}$ and have had a return $x$ so far.
We conclude from both cases that there is no optimal Markovian policy.
Figure 11: MDP serving as counterexample for the proof of the last implication. While this example has random rewards, another counterexample with random transitions and deterministic rewards can be designed, in the same way as in Remark 11.
G.5 Proof of Proposition 5
@@
We proceed in the same way as for Proposition 2 by providing a counterexample. We consider the utility function $\mathcal{U}:x\mapsto 1-\exp(-\beta x)$ with $\beta=\frac{1}{2}$, and the same $w^{+}$ function as in the proof of Proposition 2:
@@
We also set $w^{-}=0$. However, we consider another MDP. Our MDP has three states: an initial state $s_{0}$, an intermediate state $s_{1}$, and a terminal state $s_{2}$. There are two actions: A and B. All trajectories start in $s_{0}$. Any action from $s_{0}$ leads to $s_{1}$ with probability $1$ and yields reward $+1$ with probability $\frac{1}{2}$ and 00 otherwise. The action taken when in $s_{0}$ is completely irrelevant. Any action taken in $s_{1}$ leads to $s_{2}$ with probability $1$ and the episode stops as soon as $s_{2}$ is reached. When taking action A in $s_{1}$, the reward is either 00 or $+2$, with probability $\frac{1}{2}$ each. When taking action B in $s_{1}$, the reward is $+1$ with probability $1$. All policies in $\Pi_{NM}$ can be described by $(p_{\text{start}},p_{0},p_{1})$, where $p_{\text{start}}$ is the probability of choosing action A when in $s_{0}$, $p_{0}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward 00 and $p_{1}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward $1$. $p_{\text{start}}$ is irrelevant to the performance of the policy so we can ignore it. The set of Markovian policies here is the set of policies such as $p_{0}=p_{1}$. $\mathbb{C}(\pi)$ is a piecewise affine function of $p_{0}$ and $p_{1}$ and it can therefore be directly maximized. We omit the calculations here: one can check that the best achievable CPT value for Markovian policies is $\approx 0.616$ for $p_{0}=p_{1}=0.4$ but that a CPT value of $\approx 0.625$ is achievable for $p_{0}=0$ and $p_{1}=0.4$, proving the lemma.
Appendix H Proofs and Additional Details for Section 4
H.1 Proof of Theorem 6
@@
As we consider a finite horizon setting with finite state and action spaces, the integral on trajectories $\tau$ are in fact finite sums, allowing us to differentiate freely. We still write the proof with $\int$ signs, signalling our hope that, under some technical assumptions, our proof could be generalized to a setting with infinite horizon and/or infinite state and action spaces.
@@
$\mathbb{C}(X)=\int_{z=0}^{+\infty}w(\mathbb{P}(\mathcal{U}(X)>z)dz=\int_{z=0}^{+\infty}w\left(\int_{\tau\text{ such as }\mathcal{U}(R(\tau))>z}\rho_{\theta}(\tau)d\tau\right)dz\,,$ (16)
@@
$\nabla_{\theta}\mathbb{C}(X)$ $=\int_{z=0}^{+\infty}w^{\prime}(\mathbb{P}(\mathcal{U}(X>z))\nabla_{\theta}\left(\int_{\tau\text{ such as }\mathcal{U}(R(\tau))>z}\rho_{\theta}(\tau)d\tau\right)dz$ (18)
$=\int_{z=0}^{+\infty}w^{\prime}(\mathbb{P}(\mathcal{U}(X>z))\int_{\tau\text{ such as }\mathcal{U}(R(\tau))>z}\nabla_{\theta}\rho_{\theta}(\tau)d\tau dz$ (19)
[…]
@@
[…]
@@
H.2 Alternative Practical Procedure for Computing Stochastic Policy Gradients

### v3

@@ H.1 The need for stochastic policies
H.2 Proof of Proposition 2
@@
We proceed by providing a counterexample. We consider the utility function $\mathcal{U}:x\mapsto 1-\exp(-\beta x)$ with $\beta=\frac{1}{2}$, and the weight function:
@@
We also set $w^{-}=0$. However, we consider another MDP. Our MDP has three states: an initial state $s_{0}$, an intermediate state $s_{1}$, and a terminal state $s_{2}$. There are two actions: A and B. All trajectories start in $s_{0}$. Any action from $s_{0}$ leads to $s_{1}$ with probability $1$ and yields reward +1+1 with probability $\frac{1}{2}$ and 0 otherwise. The action taken when in $s_{0}$ is completely irrelevant. Any action taken in $s_{1}$ leads to $s_{2}$ with probability $1$ and the episode stops as soon as $s_{2}$ is reached. When taking action A in $s_{1}$, the reward is either 0 or +2+2, with probability $\frac{1}{2}$ each. When taking action B in $s_{1}$, the reward is +1+1 with probability $1$. All policies in $\Pi_{NM}$ can be described by $(p_{\text{start}},p_{0},p_{1})$, where $p_{\text{start}}$ is the probability of choosing action A when in $s_{0}$, $p_{0}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward 0 and $p_{1}$ is the probability of choosing action A in $s_{1}$ if the transition from $s_{0}$ to $s_{1}$ yielded reward $1$. $p_{\text{start}}$ is irrelevant to the performance of the policy so we can ignore it. The set of Markovian policies here is the set of policies such as $p_{0}=p_{1}$. $\mathbb{C}(\pi)$ is a piecewise affine function of $p_{0}$ and $p_{1}$ and it can therefore be directly maximized. We omit the calculations here: one can check that the best achievable CPT value for Markovian policies is $\approx 0.616$ for $p_{0}=p_{1}=0.4$ but that a CPT value of $\approx 0.625$ is achievable for $p_{0}=0$ and $p_{1}=0.4$, proving the lemma.
Appendix I Proofs and Additional Details for Section 3
I.1 Proof of Theorem 3
@@
As we consider a finite horizon setting with finite state and action spaces, the integral on trajectories $\tau$ are in fact finite sums, allowing us to differentiate freely. We leave the interesting and technical question of the extension to continuous state-action spaces to future work.
@@
$\mathbb{C}(X)=\int_{z=0}^{+\infty}w(\mathbb{P}(\mathcal{U}(X)>z)dz=\int_{z=0}^{+\infty}w\left(\int_{\tau\,\text{s.t.}\,\mathcal{U}(R(\tau))>z}\rho_{\theta}(\tau)d\tau\right)dz\,$, (9)
@@
$\nabla_{\theta}\mathbb{C}(X)$ $=\int_{z=0}^{+\infty}w^{\prime}(\mathbb{P}(\mathcal{U}(X>z))\nabla_{\theta}\left(\int_{\tau\,\text{s.t.}\,\mathcal{U}(R(\tau))>z}\rho_{\theta}(\tau)d\tau\right)dz$
$=\int_{z=0}^{+\infty}w^{\prime}(\mathbb{P}(\mathcal{U}(X>z))\int_{\tau\,\text{s.t.}\,\mathcal{U}(R(\tau))>z}\nabla_{\theta}\rho_{\theta}(\tau)d\tau dz$
$=\int_{\tau}\int_{z=0}^{\mathcal{U}(R(\tau))}w^{\prime}(\mathbb{P}(\mathcal{U}(X)>z))\nabla_{\theta}\rho_{\theta}(\tau)dzd\tau$
$=\int_{\tau}\phi(\mathcal{U}(R(\tau)))\nabla_{\theta}\rho_{\theta}(\tau)d\tau\,$, (11)
where $\phi(t):=\int_{z=0}^{t}w^{\prime}(\mathbb{P}(\mathcal{U}(X)>z))dz$ for any real t.t\,.
@@
$\nabla_{\theta}\mathbb{C}(X)=\int_{\tau}\phi(\mathcal{U}(R(\tau)))\rho(\tau)\nabla_{\theta}\log\rho(\tau)d\tau=\mathbb{E}_{\tau\sim\rho}[\phi(\mathcal{U}(R(\tau)))\nabla_{\theta}\log\rho(\tau)]\,.$
@@
I.2 Alternative Practical Procedure for Computing Policy Gradients